What is the 15th term of the sequence -13, -9, -5, -1, . . . ?
Answer by robertb(4012)

a1 = -13, d = 4, so an = -13 + (n-1)4 = -13 + 4n-4 = 4n - 17. Thus a15 = 4*15-17 = 60-17 = 43.